Suppose EP = -0.5.

-0.5 = %D Q / %D P

If %D P is 10, %D Q = -5, so total expenditures ­ by 5% - Why?

Suppose EP = -1.0.

-1.0 = %D Q / %D P

If %D P is 10, %D Q = -10, so total expenditures are unchanged - Why!
Answer #1

Total expenditure = TE= P*Q

Taking log both sides

LogTE = logP + logQ

Differentiating with respect to time

% change in TE = % change in P + % change in Q

1) change in total expenditure = 10-5= 5%

2) total expenditure = 10-10= 0

When demand ia unit elastic percentage change in quantity demanded is equal to percentage change in price that makes the total expenditure unchanged.